A small sphere carrying 6.5 nC

nCof charge is placed 100 mm

from a small sphere carrying 4.0 nC

of charge. Assume the spheres are tiny relative to the 100-mm

separation distance.

At what position from the particle with 6.5 nC

of charge on the line joining the spheres is the electric field magnitude zero?

Express your answer with the appropriate units.




1
Expert's answer
2021-08-31T17:32:53-0400
"\frac{6.5}{a^2}=\frac{4}{(a-0.1)^2}\\a=0.46\ m"